OpenStudy (shoxa27):

a square and a regular hexagon are placed so that they have a common side. find m

OpenStudy (shoxa27):



OpenStudy (anonymous):

do you know how to find the interior angle of a square and of a hexagon?

OpenStudy (shoxa27):

not really

OpenStudy (anonymous):

sum of interior angles = 180(n-2) where n is the number of sides. You can then find each measure of the angle by dividing by the total number of angles. so for the square you would do: 180(4-2) = 360 degrees. Now find each angle by dividing by 4, so 360/4 = 90

OpenStudy (anonymous):

Try it for the hexagon